OpenStudy (anonymous):

For the line segment whose endpoints are X (1, 2) and Y (6, 7), find the x value for the point located 1 over 3 the distance from X to Y. 3.5 3.3 <- my answer correct me if I'm wrong thank you 2.7 2.4

OpenStudy (anonymous):

Take the x-distance and y-distance, and multiply each by 1/3. \[\Delta*x=6-1=5\] \[\Delta*y=7-2=5\] 5/3=1.6666...., so the point 1/3 the distance between X and Y is (2.667,2.667)

OpenStudy (anonymous):

So \[x \approx 2.7\]

OpenStudy (anonymous):

thank you for explaining